# CounterCurate: Enhancing Physical and Semantic Visio-Linguistic Compositional Reasoning via Counterfactual Examples

### v2

@@ 4.2 Evaluating Positional Understanding
Again, to the best of our knowledge, we are the first to comprehensively evaluate a multimodal model’s ability in understanding the left-and-right and/or above-and-below positioning relations between objects. We conduct a 4:1 train-test split on Flickr30k-Positions and use the test subset as the benchmark. We further separate this dataset into 3 subsets: left-and-right only, above-and-below only, and both. We evaluate our fine-tuned CLIP models by comparing the CLIP similarity scores between the four image-caption pairs, $(C,I),(C,I^{\prime}),(C^{\prime},I),(C^{\prime},I^{\prime})$ from each group of data in the dataset, where $(C,I,C^{\prime},I^{\prime})$ denotes the positive and negative image/caption, respectively. The model receives a score of 0.5 if $s_{(C,I)}>s_{(C,I^{\prime})}$ and another score of 0.5 if $s_{(C^{\prime},I)}<s_{(C^{\prime},I^{\prime})}$, where $s_{(C,I)}$ is the CLIP cosine-similarity score between caption $C$ and image $I$. For LLaVA-1.5, we simply query the model to choose between the positive and negative captions when presented with the ground truth image. The results are shown in Table 1.
@@
As we hypothesized, LMMs are indeed largely oblivious to the objects’ positioning in the image, which is especially manifested in the vanilla CLIP’s performance, which is only marginally better than random guessing. Vanilla LLaVA-1.5 shows slightly better performance.
After fine-tuning with the training split of Flickr30k-Positions, both models perform significantly better across all subsets. Specifically, for the mixed case, CLIP improves by 33%, and LLaVA achieves a high accuracy of 96%. These results demonstrate that CounterCurate is highly effective across different kinds of multimodal models.
@@ 4.4 Evaluating Semantic Compositional Reasoning
For example, CounterCurate fine-tuned CLIP surpasses NegCLIP on average as well as in two main categories. Note that the source of fine-tuning data, Flickr30k-Entities, contains much fewer image-caption pairs than COCO-Captions (around 100k image-caption pairs) where NegCLIP is trained. Furthermore, when prompting GPT-4V to generate the negative captions, we intentionally prompted GPT-4V to produce None when there is no feasible condition to conduct “swapping". Such two factors lead to the fact that our data curation pipeline results much fewer negative samples for the “swap” category, which is around 2.5k. We argue that using our pipeline on other datasets that can generate more of the “swap” attribute shall lead to larger improvements in performance.
LLaVA-1.5 also performed better in all three categories after fine-tuning. It is also surprising that our fine-tuned model outperforms the SOTA LMM GPT-4V both on average and in two of the categories, most significantly over the “add” category. We observe improvements on other datasets Diwan et al. (2022) as well in Appendix E.
CounterCurate shows better performance compared to the prior rule-based methods or less desirable models that generates negatives. These significant improvements show that fine-tuning with accurate and hard negative samples is important, again demonstrating the effectiveness of our data curation and fine-tuning pipeline for both contrastive models and text generation models for improving semantic counterfactual understanding.
@@ 4.5 In-Depth Analysis
As shown in Table 4, even though using either the negative caption or the negative image can improve performance, the scores are significantly worse than when using both elements to fine-tune. This demonstrates that CounterCurate, which incorporates both negative captions and negative images, is necessary to achieve desirable improvements. Grouping also delivers further improvements compared to not using this strategy.
@@
Performance on zero-shot vision-language tasks
To evaluate whether fine-tuning on the counterfactual data hurts the original zero-shot vision-language performance, we compare the image and text retrieval performance on MSCOCO Lin et al. (2015) of the original CLIP model and CounterCurate fine-tuned model on Flickr30k-Attributes. The results in Table 5 show that on average, the model fine-tuned via CounterCurate maintains performance with marginal difference compared to the original CLIP model.
@@ Acknowledgement
We acknowledge that DALLE-3, one of the generative models we use, is a closed-source text-to-image generation model, which makes it difficult to systematically analyze its behavior.

### v3

@@ 4.2 Evaluating Positional Understanding
To the best of our knowledge, we are the first to comprehensively evaluate a multimodal model’s ability in understanding the left-and-right and/or above-and-below positioning relations between objects. We conduct a 4:1 train-test split on Flickr30k-Positions and use the test subset as the benchmark. We further separate this dataset into 3 subsets: left-and-right only, above-and-below only, and both. We evaluate our fine-tuned CLIP models by comparing the CLIP similarity scores between the four image-caption pairs, $(C,I),(C,I^{\prime}),(C^{\prime},I),(C^{\prime},I^{\prime})$ from each group of data in the dataset, where $(C,I,C^{\prime},I^{\prime})$ denotes the positive and negative image/caption, respectively. The model receives a score of 0.5 if $s_{(C,I)}>s_{(C^{\prime},I)}$ and another score of 0.5 if $s_{(C,I^{\prime})}<s_{(C^{\prime},I^{\prime})}$, where $s_{(C,I)}$ is the CLIP cosine-similarity score between caption $C$ and image $I$. For LLaVA-1.5, we simply query the model to choose between the positive and negative captions when presented with the ground truth image. The results are shown in Table 1.
@@
As we hypothesized, LMMs are indeed largely oblivious to the objects’ positioning in the image, which is especially manifested in the vanilla CLIP’s performance, which is only marginally better than random guessing. Vanilla LLaVA-1.5 shows only slightly better performance.
After fine-tuning with the training split of Flickr30k-Positions, both models perform significantly better across all subsets. Specifically, for the mixed case, CLIP improves by 33% points, and LLaVA achieves a high accuracy of 96%. These results demonstrate that CounterCurate is highly effective across different kinds of multimodal models.
@@ 4.4 Evaluating Semantic Compositional Reasoning
For example, CounterCurate fine-tuned CLIP surpasses NegCLIP on average as well as in two main categories. Note that the source of fine-tuning data, Flickr30k-Entities, contains much fewer image-caption pairs than COCO-Captions (around 100k image-caption pairs) where NegCLIP is trained. Furthermore, when prompting GPT-4V to generate the negative captions, we intentionally prompt GPT-4V to produce None when there is no feasible condition to conduct “swapping". These two factors lead to our data curation pipeline resulting in much fewer negative samples for the “swap” category, which is around 2.5k. We argue that using our pipeline on other datasets that can generate more of the “swap” attribute should lead to larger improvements in performance.
LLaVA-1.5 also performs better in all three categories after fine-tuning. It is also surprising that our fine-tuned model outperforms the SoTA LMM GPT-4V both on average and in two of the categories, most significantly for the “add” category. We observe improvements on other datasets Diwan et al. (2022) as well in Appendix E. The promising results of LLaVA, such as the score of 94.17% on SugarCrepe surpassing GPT-4V and a 10% improvement on Flickr30k-Position’s above-below subset compared to GPT-4V, showcases that CounterCurate effectively improves LLaVA’s semantic and physical reasoning capabilities, and we thus identify CounterCurate as a valid candidate towards improving SoTA LMMs such as GPT-4V as well.
CounterCurate shows better performance compared to the prior rule-based methods or less desirable models that generates negatives. These significant improvements show that fine-tuning with accurate and hard negative samples is important, again demonstrating the effectiveness of our data curation and fine-tuning pipeline for both contrastive models and generative models for improving semantic counterfactual understanding.
@@ 4.5 In-Depth Analysis
As shown in Table 4, even though using either the negative caption or the negative image can improve performance, the scores are significantly improved when using both elements to fine-tune. This demonstrates that CounterCurate, which incorporates both negative captions and negative images, is necessary to achieve desirable improvements. Grouping also delivers further improvements compared to not using this strategy.
@@
MMBench MMBench_cn LLaVA_W POPE ScienceQA MM-Vet VizWiz MME
LLaVA-1.5 67.7 63.6 70.7 85.9 71.6 35.4 53.6 1531
+ CounterCurate 68.6 64.1 71.1 85.6 72.1 39.5 54.2 1510
Table 5: Evaluating our fine-tuned LLaVA-1.5 on its original benchmarks.
@@
Performance on zero-shot vision-language tasks
We evaluate whether fine-tuning on the counterfactual data hurts the original zero-shot vision-language performance. For CLIP, we compare the image and text retrieval performance on MSCOCO Lin et al. (2014) of the original CLIP model and CounterCurate fine-tuned model on Flickr30k-Attributes. The results in Table 6 show that on average, the model fine-tuned via CounterCurate maintains performance with marginal difference compared to the original CLIP model. In Appendix I, we also show that fine-tuning with CounterCurate significantly improves CLIP’s performance on several common downstream tasks. For LLaVA, we mix Flickr30k-Attributes with LLaVA’s original training data before fine-tuning our model. Compared with vanilla LLaVA-1.5 in Table 5 on the same benchmarks from Liu et al. (2023a), the CounterCurate-finetuned LLaVA overall performs similarly (or even slightly better). From these studies, we see that improving LMMs’ semantical compositional reasoning with Flickr30k-Attributes does not hurt downstream performance of both contrastive and generative models.
@@ Acknowledgement
We acknowledge that DALLE-3 and GPT4-V are closed-source models, which makes it difficult to systematically analyze their behavior.
@@ Appendix G More Results on SugarCrepe
Appendix H Ratio of Augmentation
We study the level of effectiveness our dataset Flickr30k-Attributes brings upon the models if we choose to reduce the amount of curated data used during fine-tuning. As we observe in Table 10 and especially Table 11, using 100% of our dataset yields the greatest amount of improvements in LLaVA and CLIP’s semantical compositional reasoning capabilities. This prompt us to further curate data using the same method in the future that shall bring forth even more competent models.
Type LLaVA-1.5 + CounterCurate (50%) (25%) (10%) (5%) (2.5%)
add_att 80.64 95.09 94.22 90.90 88.58 88.44 86.27
add_obj 87.83 97.82 97.58 95.44 94.67 93.79 90.45
replace_att 92.64 93.53 93.27 91.50 92.51 91.62 92.64
replace_obj 96.73 98.37 98.37 98.24 98.06 97.88 97.28
replace_rel 87.13 85.92 84.92 85.63 86.77 88.12 88.19
swap_att 86.34 92.79 90.54 89.19 88.44 90.09 89.79
swap_obj 84.90 85.72 83.67 81.63 82.04 84.08 86.12
add_average 86.02 97.13 96.73 94.29 93.13 92.44 89.39
replace_average 92.38 92.83 92.40 92.24 92.79 93.02 93.00
swap_average 85.95 90.89 88.69 87.15 86.71 88.47 88.80
Overall 89.27 94.17 93.54 92.38 92.18 92.26 91.17
Table 10: Ratio of augmentation ablation study on fine-tuning LLaVA-1.5.
Type CLIP + CounterCurate (50%) (25%) (10%) (5%) (2.5%)
add_att 87.14 90.34 89.23 88.11 86.85 87.09 87.05
add_obj 77.89 86.70 81.21 79.91 78.46 78.03 77.74
replace_att 82.61 87.94 84.77 82.99 83.37 82.86 82.61
replace_obj 93.76 95.94 94.79 94.37 93.88 93.76 93.76
replace_rel 68.91 76.24 72.97 70.19 69.55 69.34 69.13
swap_att 67.41 73.57 68.76 66.81 66.66 66.51 67.26
swap_obj 60.00 68.57 62.85 62.85 63.26 62.44 60.40
add_average 80.21 87.62 83.22 81.97 80.57 80.31 80.08
replace_average 82.39 87.10 84.76 83.20 82.83 82.60 82.47
swap_average 65.42 72.22 67.17 65.75 65.75 65.42 65.42
Overall 79.54 85.49 85.65 82.07 80.64 79.94 79.68
Table 11: Ratio of augmentation ablation study on fine-tuning CLIP.
Appendix I More evaluation on CLIP downstream tasks
In Table 12, we compare vanilla CLIP with CounterCurate-finetuned CLIP on Flickr30k-Attributes on some common downstream tasks. The results show that CounterCurate generally improves performance.
Benchmark CounterCurate LAION-2b
ImageNet-a Hendrycks et al. (2021) 26.8 26.25
kitti-closest-vehicle-distance Ginger et al. (2021) 30.1 26.16
eurosat Helber et al. (2019) 51.59 48.13
voc2007_multilabel Everingham et al. (2015) 82.49 79.63
dmlab Zhai et al. (2019) 16.81 15.97
pcam Veeling et al. (2018) 60.58 59.78
fgvc_aircraft Maji et al. (2013) 25.26 24.54
Table 12: Comparing our fine-tuned CLIP with LAION-2b pretrained CLIP on common downstream tasks.